### v1

# On the Generalization Error of Differentially Private Algorithms Via Typicality

Abstract
We study the generalization error of stochastic learning algorithms from an information-theoretic perspective, with a particular emphasis on deriving sharper bounds for differentially private algorithms. It is well known that the generalization error of stochastic learning algorithms can be bounded in terms of mutual information and maximal leakage, yielding in-expectation and high-probability guarantees, respectively. In this work, we further upper bound mutual information and maximal leakage by explicit, easily computable formulas, using typicality-based arguments and exploiting the stability properties of private algorithms. In the first part of the paper, we strictly improve the mutual-information bounds by Rodríguez-Gálvez et al. (IEEE Trans. Inf. Theory, 2021). In the second part, we derive new upper bounds on the maximal leakage of learning algorithms. In both cases, the resulting bounds on information measures translate directly into generalization error guarantees.
I Introduction
A randomized statistical learning algorithm $P_{W|S}$ is defined as a probabilistic mapping from a training dataset $S:=\left(Z_{1},\ldots,Z_{N}\right)\in\mathcal{Z}^{N}$ consisting of $N$ independent and identically distributed (i.i.d.) data points from distribution $P_{Z}$ (which is unknown) on the space $\mathcal{Z}$ to a hypothesis $W$ from the hypothesis space $\mathcal{W}$. We assume $\mathcal{Z}$ is finite and $\mathcal{W}$ is discrete, like [1]. The generalization performance of this learning algorithm is measured by the difference between the expected loss $\mathbf{E}_{P_{Z}}\left[\ell(W,Z)\right]$ and the empirical loss $\frac{1}{N}\sum^{N}_{i=1}\ell(W,Z_{i})$, where $\ell(\cdot,\cdot)$ is the loss function. We assume $\ell(\cdot,\cdot)$ is $\sigma$-sub-Gaussian in this paper, i.e., for all $\lambda\in\mathbb{R}$ and all $w\in\mathcal{W}$,
$\mathbf{E}_{Z\sim P_{Z}}\big[\exp\big(\lambda(\mathbf{E}[\ell(w,Z)]-\ell(w,Z))\big)\big]\leq\exp\big(\lambda^{2}\sigma^{2}/2\big).$
One of the key challenges in statistical learning theory is to guarantee that a learning algorithm can “generalize well”, i.e., the generalization error defined as
$\mathrm{gen}(W,S):=\mathbf{E}_{P_{Z}}\left[\ell(W,Z)\right]-\frac{1}{N}\sum^{N}_{i=1}\ell(W,Z_{i}),$ (1)
is bounded and hence how much the algorithm overfits to the data is under control.
A vast array of studies have been proposed for this objective. To guarantee that the trained algorithm can generalize well, the theory of uniform convergence [2] states that the output has to be “sufficiently simple”. Following this idea, bounds in terms of the VC-dimension and the Rademacher complexity have been proposed [3, 4]. However, they make no reference to the algorithm. To make a specific algorithm generalize, various types of methods have been studied, e.g., based on compression schemes [5] or uniform stability [6]. Motivated by controlling the bias in data analysis using the mutual information between the empirical risks of the hypotheses and the algorithm output [7], it was shown in [8] that the expectation of the generalization error can be bounded by the mutual information $I(S;W)$ between the training dataset $S$ and the output hypothesis $W$:
$\mathbf{E}\left[\big|\mathrm{gen}(W,S)\big|\right]\leq\sqrt{\frac{2\sigma^{2}}{N}I(S;W)}.$ (2)
This provides the intuition that “an algorithm that leaks little input information generalizes well” [9, 10], which partially coincides with the idea of private algorithms. There exist various definitions of privacy, including total variation distance [11], differential privacy [12, 13], and max-information [14]. See Section II-B for a detailed discussion.
Maximal leakage $\mathcal{L}(S\rightarrow W)$ [15] of an algorithm $P_{W|S}$, as another information leakage measure, has been proposed and discussed recently [16, 17, 18, 19]. In [18], it has been shown that the generalization error can also be bounded by $\mathcal{L}(S\rightarrow W)$: given $\eta\in(0,1)$,
$\mathbf{P}\left(\big|\mathrm{gen}(W,S)\big|\geq\eta\right)\leq 2\exp\left(\mathcal{L}(S\rightarrow W)-\frac{N\eta^{2}}{2\sigma^{2}}\right).$ (3)
One major meaningful difference between (3) and (2) is that (3) is a high-probability (a.k.a. single-draw [20, 21]) bound that admits an exponentially decaying (in the size of the training dataset $N$) bound on the probability of having a large generalization error, which is generally untrue for mutual-information-based bounds that only guarantee a linearly decaying bound [10].
For generalization error bounds in terms of other information measures, see [22, 23, 24, 18, 21, 25, 26]. The connection between privacy measures and generalization has been widely discussed in [14, 27, 28, 18, 1, 29]. Note that privacy is tightly related to compression [30, 31], and compressibility can be used to bound the generalization error as well [5, 23, 24].
In [1], an approach based on typicality [32] has been used to bound the generalization error of private algorithms, and their upper bounds are in more explicit forms, since mutual information is sometimes hard to calculate.
In this paper, our contributions are two-fold. In the first part of this paper (Section III), we utilize the stability property of differentially private algorithms to further strengthen the mutual information bounds in [1], and show that our results strictly improve the existing bounds. In the second part of this paper (Section IV), we propose novel upper bounds on the maximal leakage, which directly translate to upper bounds on mutual information as well [15].
Notations
We use calligraphic letters (e.g., $\mathcal{W}$), capital letters (e.g., $W$) and lower-case letters (e.g., $w$) and to denote sets, random variables and instances, respectively. We assume all random variables are from finite alphabets and logarithm and entropy are to the base $e$ unless otherwise stated. For a statement $S$, $\mathds{1}_{\{S\}}$ is its indicator, i.e., $\mathds{1}_{\{S\}}$ is $1$ if $S$ holds, and 0 otherwise.
II Preliminaries
We use $S:=\left(Z_{1},\ldots,Z_{N}\right)$ to denote a training dataset, which is a collection of $N$ instances $Z_{i}\in\mathcal{Z}$ i.i.d. from $P_{Z}$. A stochastic learning algorithm $\mathcal{A}:\mathcal{Z}^{N}\rightarrow\mathcal{W}$ can be characterized by the conditional distribution $P_{W|S}$, i.e., $S$ is a random variable with distribution $P_{S}=P_{Z}^{\otimes N}$ over $\mathcal{Z}^{N}$, where $|\mathcal{Z}|$ is finite. We consider algorithms that are permutation-invariant: $\mathcal{A}$ operates on a set instead of an ordered sequence; the hypothesis generated by $\mathcal{A}$ with dataset $S$ is equal in distribution to the hypothesis generated by $\mathcal{A}$ with a permuted dataset $\mathcal{P}(S)$, i.e., $\mathcal{A}(S)\stackrel{{\scriptstyle d}}{{=}}\mathcal{A}(\mathcal{P}(S))$ for all permutations $\mathcal{P}$. This is a common assumption in the generalization analysis of learning theory; see also [6, 33].
In this paper, we upper-bound the generalization error (1) via either an in-expectation bound of the form (2) or a high-probability bound of the form (3). There are other types of generalization error bounds, e.g., PAC-Bayesian bounds [21, 34, 35], which may also be of interest in the future. We consider $\sigma$-sub-Gaussian loss functions in this paper; note that if the loss function $\ell$ is bounded on $[a,b]$, then it is (ba)/2(b-a)/2-sub-Gaussian.
II-A Method of Types
Consider $Z^{N}:=(Z_{1},\ldots,Z_{N})$ where $Z_{i}\stackrel{{\scriptstyle\mathrm{i.i.d}}}{{\sim}}P_{Z}$ for $i=1,\ldots,N$. We can then define the type [32] of $Z^{N}$ as follows.
Definition 1 ([32]).
The type of a sequence $Z^{N}\in\mathcal{Z}^{N}$ is the relative proportion of occurrences of each symbol of $\mathcal{Z}$, i.e., $T_{Z^{N}}(t)=\mathsf{N}(t\,|\,Z^{N})/N$ for all $t\in\mathcal{Z}$, where $\mathsf{N}(t\,|\,Z^{N})$ is the number of times the symbol $t$ occurs in the sequence $Z^{N}$. The set of all possible types of sequences of elements from $\mathcal{Z}$ of length $N$ is denoted by $\mathcal{T}_{\mathcal{Z},N}$.
The type and its variants were used to prove achievabilities of network information theory problems, either asymptotic [36] or nonasymptotic [37, 38]. For the latter, also see other methods [39, 40, 41]. Even though the number of items in $\mathcal{Z}^{N}$ scales exponentially in $N$, it is known that the total number of types only scales polynomially: [32] gives $|\mathcal{T}_{\mathcal{Z},N}|\leq(N+1)^{|\mathcal{Z}|}$, which was marginally improved by [1]:
Claim 1 (Claim 1 of [1]).
$|\mathcal{T}_{\mathcal{Z},N}|\,\leq\,(N+1)^{|\mathcal{Z}|-1}$, with equality if and only if $|\mathcal{Z}|=2$.
Moreover, the property of the type that will be utilized in this paper is that it is equal to that of any permutation of the same sequence, i.e., $T_{Z^{N}}=T_{\mathcal{P}(Z^{N})}$ for all permutation $\mathcal{P}$, and hence $T_{Z^{N}}$ can uniquely define a dataset $S=\{Z_{1},\ldots,Z_{N}\}$. Hence [1] can define the type of the dataset $S$ as
$T_{S}\triangleq T_{\tilde{Z}^{N}}$
where $\tilde{Z}^{N}$ is a sequence generated by all the $N$ instances of $S$, and therefore the distance between two datasets $S,S^{\prime}$ can be defined via their types $T_{S},T_{S^{\prime}}$.
Definition 2 (Definition 5 of [1]).
The distance between two datasets $S,S^{\prime}$ is the minimum number of instances that needs to be changed in $S$ to obtain $S^{\prime}$.
$\mathsf{d}(S,S^{\prime})$ $\triangleq\frac{1}{2}\sum_{a\in\mathcal{Z}}\big|\mathsf{N}(a|S)-\mathsf{N}(a|S^{\prime})\big|$
$=\frac{N}{2}\big\|T_{S}-T_{S^{\prime}}\big\|_{1}$
where $\mathsf{N}(a|S)$ denotes the number of times the symbol $a$ occurs in the dataset $S$.
Due to the permutation invariance of the assumed algorithms, the use of types is justified by the bijection between types and datasets.
II-B Differential Privacy
Differential privacy [12, 13], as one of the most celebrated privatization schemes, has been widely studied in the past years (also see its variants [42, 43, 44]):
Definition 3.
Given an algorithm $\mathcal{A}$ which induces the conditional distribution $P_{W|S}$, we say it satisfies $(\varepsilon,\delta)$-DP if for any neighboring (differ in a single data point) $(s,s^{\prime})$ and $\mathcal{Z}\subseteq\mathcal{W}$, it holds that for all measurable $W$,
$\mathbf{P}\big(W\in\mathcal{Z}|S=s\big)\leq e^{\varepsilon}\,\mathbf{P}\big(W\in\mathcal{Z}|S=s^{\prime}\big)+\delta.$
In particular, when $\delta=0$ we simply refer to it as $\varepsilon$-DP.
Differentially private algorithms are stable [28], i.e., they produce similar outputs on similar input datasets. For an $\varepsilon$-DP algorithm $\mathcal{A}$, the maximum log-likelihood ratio between any two neighboring input datasets is bounded by $\varepsilon$, and the notion of neighboring datasets relates to the distance between datasets. We say that two datasets are neighbors if they have distance $1$ (in terms of Definition 2).
One useful property is group privacy [45, Theorem 2.2]: for an $\varepsilon$-DP algorithm $\mathcal{A}$, if two (fixed) datasets $S$ and $S^{\prime}$ are at distance $k$, then the KL divergence of their output distributions is bounded by
$D_{\mathrm{KL}}\big(\mathcal{A}(S)\,\|\,\mathcal{A}(S^{\prime})\big)\leq k\varepsilon\tanh\left(k\varepsilon/2\right)\leq k\varepsilon.$ (4)
Moreover, for an $\varepsilon$-DP mechanism, in nats, we have [46]
$I(S;W)\leq\min\{\varepsilon,\varepsilon^{2}\}.$
Discussions on maximal leakage can be found in Section IV.
We are now ready to present our main results.
III Mutual Information Bounds
In [1], upper bounds on the generalization error of discrete algorithms have been proposed by upper bounding $I(W;S)$, or more precisely upper bounding the relative entropy $D_{\mathrm{KL}}(P_{W|S}\,\|\,Q_{W})$, since by the golden formula [47, Section 4.1],
$I(W;S)\leq\mathbf{E}_{s\sim P_{S}}\left[D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\right],$ (5)
where the equality is achieved if and only if $Q_{W}=P_{W}$, and the expression $D_{\mathrm{KL}}(P_{W|S}\,\|\,Q_{W})$ also appears in other generalization error characterizations, e.g., [21]. By (2), these upper bounds directly translate to generalization error bounds.
Then based on the typicality result in Claim 1, [1] has the following results:
Proposition 1 ([1]).
Let $S\in\mathcal{S}$ be a dataset of $N$ instances $Z^{N}\in\mathcal{Z}^{N}$ sampled i.i.d. from $P_{Z}$. Let also $W\in\mathcal{W}$ be a hypothesis obtained with an algorithm $\mathcal{A}$, characterized by $P_{W|S}$. For all $s\in\mathcal{S}$, there exists a distribution $Q_{W}$ such that
  • If $\mathcal{A}$ is generally discrete, then
    $D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\leq\left(|\mathcal{Z}|-1\right)\log(N+1).$ (6)
  • If $\mathcal{A}$ is also $\varepsilon$-DP and $\varepsilon\leq 1$, then
    $D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\leq(|\mathcal{Z}|-1)\log(1+e\varepsilon N).$
Combining with the in-expectation generalization error bound (2) and the golden formula (5), from (6) one can derive
$\big|\mathbf{E}_{P_{WS}}\left[\mathrm{gen}(W;S)\right]\big|\leq\sqrt{\frac{2\sigma^{2}}{N}\left(|\mathcal{Z}|-1\right)\log(N+1)}.$
In this paper, we first show a slightly improved version of Claim 1, since the proof of Claim 1 (see [1, Appendix G.1]) used a the very loose step $1+N/i\leq 1+N$ for every $i$.
Claim 2.
$|\mathcal{T}_{\mathcal{Z},N}|\,\leq\,\frac{1}{\sqrt{2\pi(|\mathcal{Z}|-1)}}\left(\frac{e}{|\mathcal{Z}|-1}\left(N+\frac{|\mathcal{Z}|}{2}\right)\right)^{|\mathcal{Z}|-1}$.
The proof of Claim 2 can be found in Appendix A.
Our Claim 2 is strictly better than Claim 1 by a constant factor <1<1, and this constant factor $(k+1)/2^{k}$ is actually an optimal factor $c$ such that $\binom{N+k}{k}\leq c(N+1)^{k}$ holds for all $N\geq 1$. Recall $k=|\mathcal{Z}|-1$, it can be dramatically tighter as $|\mathcal{Z}|$ grows. However, we note that the advantages of our theorems do not mostly come from here, we are just using Claim 2 to make our bounds as tight as possible. In the bounds presented later, we will show them utilizing both claims separately.
Lemma 2 (Lemma 2 of [1]).
Let $P$ and $Q$ be two probability distributions such that $P\ll Q$. Let $Q$ be a finite mixture of probability distributions such that $Q=\sum_{b}\omega_{b}Q_{b}$, where $\sum_{b}\omega_{b}=1$, and $P\ll Q_{b}$ for all $b$. Then, the following inequalities hold:
$D_{\mathrm{KL}}(P\|Q)$ $\leq-\log\left(\sum\nolimits_{b}\omega_{b}\exp\big({-}D_{\mathrm{KL}}(P\|Q_{b})\big)\right)$ (7)
$\leq\min_{b}\big\{D_{\mathrm{KL}}(P\|Q_{b})-\log(\omega_{b})\big\}.$ (8)
Inequalities in above lemma can already be tight and cannot be uniformly improved. However, for the application of differentially private algorithms, we can sharpen it as follows.
Lemma 3.
Let $Q$ be a mixture of $M$ probability distributions $(M<\infty)$ such that
$Q=\sum_{b=1}^{M}\omega_{b}Q_{b},\qquad\omega_{b}>0,\ \ \sum_{b=1}^{M}\omega_{b}=1.$
Fix an index $i\in\{1,\dots,M\}$. For each $b$, let $\alpha_{b,i}\in[0,1]$ be the largest constant such that for every measurable set $E$,
$Q_{b}(E)\geq\alpha_{b,i}\,Q_{i}(E).$
Then for any $P\ll Q_{i}$,
$D_{\mathrm{KL}}(P\|Q)\leq D_{\mathrm{KL}}(P\|Q_{i})-\log\left(\sum_{b=1}^{M}\omega_{b}\alpha_{b,i}\right).$
The proof of Lemma 3 can be found in Appendix B.
Remark 1.
If $Q_{b}=P_{W|S=s_{b}}$ are output distributions of an $\varepsilon$-DP algorithm and $\mathsf{d}(\cdot,\cdot)$ is the dataset distance, then group privacy implies for all measurable $E$,
$Q_{b}(E)$ $=P_{W|S=s_{b}}(E)$
$\geq e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\,P_{W|S=s_{i}}(E)$
$=e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\,Q_{i}(E).$
Hence the lemma applies with $\alpha_{b,i}=e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$ for any $b$.
Remark 2.
The bound by Lemma 3 is always at least as good as (8) since $\sum^{M}_{b=1}\omega_{b}\alpha_{b,i}\geq\omega_{i}\alpha_{i,i}=\omega_{i}$, and is strictly better whenever $\sum_{b\neq i}\omega_{b}\alpha_{b,i}>0$.
We now consider differentially private algorithms and show how to further improve the results.
We first define what do we mean by representative sets, which is from the construction presented in [1, Appendix C]. The idea is to utilize the property that a private algorithm produces statistically similar outputs for neighboring datasets. This will also be utilized in our proof.
Note all types $T_{s}$ lie inside the unit hypercube $0,1]^{|\mathcal{Z}|-1}$, and can be uniquely identified by the preceding $(|\mathcal{Z}|-1)$ dimensions. Rather than discussing types, we study the vector of counts $\mathsf{N}_{s}$ by $\mathsf{N}_{s}(a)=\mathsf{N}(a|s)$ for all $a\in\mathcal{Z}$, and hence $\mathsf{N}_{s}$ lies in a $0,N]^{|\mathcal{Z}|-1}$ hypercube. For the interval $[0,N]$, each dimension can be split to $t$ parts where integer $t\in[1:N]$, hence resulting in a $0,N]^{|\mathcal{Z}|-1}$ hypercube covering of $t^{|\mathcal{Z}|-1}$ hypercube smaller hypercubes, and the side of each small hypercube has length $l=N/t$ and $l^{\prime}=\lfloor l\rfloor+1$ atoms.
The representative sets is constructed by:
  • If $l^{\prime}$ is odd, choose the central atom as the corresponding coordinate for the center of the small hypercube.
  • If $l^{\prime}$ is even, enlarge the small hypercube in one unit and choose the center of this enlarged hypercube.
Theorem 4.
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$, with an induced distribution $P_{W|S}$ where $W\in\mathcal{W}$ is a hypothesis obtained with an algorithm $\mathcal{A}$. For all dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}$ on $\mathcal{W}$ such that for every $t\in\{1,2,\dots,N\}$,
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}\bigr)$
$\leq(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+(|\mathcal{Z}|-1)\log t-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon N}\bigr).$ (9)
Proof sketch.
Construct representative set $\mathcal{S}_{0}=\{s_{1},\dots,s_{M}\}$. Take $m:=|\mathcal{Z}|$ and $Q_{W}^{(t)}=\frac{1}{M}\sum_{b=1}^{M}P_{W|S=s_{b}}$, covering arguments give an index $i$ such that $\mathsf{d}(s,s_{i})\leq(m-1)\frac{N}{t}$. By $\varepsilon$-DP and group privacy, $P_{W|S=s_{b}}(E)\geq e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}P_{W|S=s_{i}}(E)$ for all $E$; Lemma 3 upper bounds $D_{\mathrm{KL}}(P_{W|S=s}\|Q_{W}^{(t)})$ by
$D_{\mathrm{KL}}(P_{W|S=s}\|P_{W|S=s_{i}})-\log\!\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\Big).$
For the first term, it is bounded by $\varepsilon\mathsf{d}(s,s_{i})\leq(m-1)\varepsilon N/t$.
For the second term, it vanishes when $t=1$, and when $t\geq 2$ we know $M\geq 2$, picking $j\neq i$ gives $\sum_{b}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\geq 1+e^{-\varepsilon N}$. Hence it is at most $\log M-\log(1+e^{-\varepsilon N})\leq(m-1)\log t-\log(1+e^{-\varepsilon N})$. The proof is complete by combining these arguments, and we can minimize over $t$ as well. ∎
The proof of Theorem 4 can be found in Appendix C. It is easy to see that Theorem 4 strictly improves [1, Proposition 2].
In [1, Proposition 3], a different idea of counting only the necessary hypercubes was employed to further sharpen the bounds. Here, we adopt the similar idea and combine it with our technique to further strengthen our bounds.
Theorem 5.
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$. Fix $|\mathcal{Z}|\geq 2$ and $N\geq 1$. For every $t\in\{1,2,\dots,N\}$ and every dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}^{(t)}$ such that
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}^{(t)}\bigr)$
$\leq(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+(|\mathcal{Z}|-1)\log\left(t+\frac{|\mathcal{Z}|-2}{2}\right)$
$\qquad-\log((|\mathcal{Z}|-1)!)-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr).$ (10)
The proof of Theorem 5 can be found in Appendix D.
Furthermore, similar to the idea of [1, Proposition 4] by employing the strong typicality, we can further improve our bounds and derive the following result.
Theorem 6.
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$. Fix $|\mathcal{Z}|\geq 2$ and $N\geq 1$. For every $t\in\{1,2,\dots,\lfloor 2\sqrt{N\log N}\rfloor\}$ and every dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}^{(t)}$ such that
$I(S;W)\leq\frac{\varepsilon|\mathcal{Z}|\sqrt{N\log N}}{t}-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr)$
$\quad\,\,\,\,\,\,+\frac{2|\mathcal{Z}|\varepsilon}{N}+\mathds{1}_{\{t\geq 2\}}\min\{|\mathcal{Z}|\log t,\log(|\mathcal{Z}|t^{|\mathcal{Z}|-1})\}$ (11)
where $\Delta_{t}:=\lceil 2\sqrt{N\log N}/t\rceil+1$.
The proof of Theorem 6 can be found in Appendix E.
Note (11) can be written in an equivalent form
$I(S;W)\leq\frac{\varepsilon|\mathcal{Z}|\sqrt{N\log N}}{t}-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr)$
$\quad\,\,\,\,\,\,+\frac{2|\mathcal{Z}|\varepsilon}{N}+\mathds{1}_{\{t\geq 2\}}\cdot\left(|\mathcal{Z}|\log t-\left[\log\left(\frac{t}{|\mathcal{Z}|}\right)\right]_{+}\right)$
by considering $\min\{x,y\}=x-[x-y]_{+}$ where $u]_{+}:=\max\{u,0\}$. We choose to present (11) in Theorem 6 since the meaning of each term is clearer.
We plot two representative scenarios and show a comparison between our bounds and the bounds in [1].
  1. 1.
    Figure 1: a small dataset with $N=1\times 10^{3}$ and $|\mathcal{Z}|=2$. The setting is typical in applications with limited sample sizes and binary outcomes/features, such as clinical studies or small A/B tests, where we collect data from thousands of patients/users and each record carries a binary attribute (e.g., responder/non-responder, click/no-click).
  2. 2.
    Figure 2: a large dataset with $N=1\times 10^{7}$ and $|\mathcal{Z}|=1\times 10^{6}$. This regime arises in large-scale telemetry or interaction logs with high-cardinality categorical values, such as search queries, URLs, item IDs, or token IDs, where one observes tens of millions of events drawn from a vocabulary on the order of millions.
Figure 1: Dataset is of size $1\times 10^{3}$ with $|\mathcal{Z}|=2$.
Figure 2: Dataset is of size $1\times 10^{7}$ with $|\mathcal{Z}|=1\times 10^{6}$.
IV Maximal Leakage Bounds
In this section we consider the bound on maximal leakage, also by a typicality-based approach. By (3), these upper bounds directly translate to generalization error bounds.
Definition 4.
Given a joint distribution $P_{SW}$ on finite alphabets $\mathcal{S},\mathcal{W}$, we say the maximal leakage from $S$ to $W$ is
$\mathcal{L}(S\to W)=\log\sum_{w\in\mathcal{W}}\max_{s\in\mathcal{S}:P_{X}(s)>0}P_{W|S}(w|s).$
Note $\mathcal{L}(S\to W)$ is the Sibson’s mutual information [48, 49] of order infinity, and it upper bounds mutual information. For an $\varepsilon$-DP mechanism, in nats, we have [15]
$\mathcal{L}(S\rightarrow W)\leq\min\{N\varepsilon,\log|\mathcal{W}|\}.$
The key advantage of maximal leakage is that it depends on the sample distribution only through its support. Consequently, it is independent of the precise distribution over the samples and is particularly amenable to analysis, especially in additive noise settings [18, 29].
Note that the usual way to keep the maximal leakage of an algorithm bounded is to add noise (e.g., Laplacian or Gaussian [18]), which aligns with common differential privacy mechanisms [12]. Hence, for the study of the generalization error of differentially private algorithms, compared to bounding mutual information, it is more natural to bound the maximal leakage.
We introduce a simple bound that will be useful.
Lemma 7.
Let $P_{1},\ldots,P_{M}$ be probability mass functions on a countable set $\mathcal{W}$. Then
$\sum_{w\in\mathcal{W}}\max_{1\leq i\leq M}P_{i}(w)\leq M.$
Moreover, equality holds if and only if for every $w\in\mathcal{W}$, there exists at most one index $i$ such that $P_{i}(w)>0$.
The proof of Lemma 7 can be found in Appendix F.
IV-A General Algorithms
We first provide a general bound on maximal leakage, where $\mathcal{A}=P_{W|S}$ is not necessarily differentially private.
Theorem 8.
Let $S=(Z_{1},\ldots Z_{N})\in\mathcal{Z}^{N}$ be a dataset of $N$ instances, each sampled i.i.d. from $P_{Z}$. Let also $W\in\mathcal{W}$ be a hypothesis obtained with an algorithm $\mathcal{A}$, characterized by $P_{W|S}$. Then the maximal leakage from the dataset to the output hypothesis satisfies
$\mathcal{L}(S\rightarrow W)\leq\left(|\mathcal{Z}|-1\right)\log(N+1).$
The proof of Theorem 8 can be found in Appendix G.
Note above result directly recovers [1, Proposition 1] (also see (6)), considering maximal leakage upper-bounds mutual information. Hence we provide an alternative proof.
If we utilize the typicality bound in our Claim 2 instead, above Theorem 8 can be strengthened to
$\mathcal{L}(S\rightarrow W)$
$\leq\log\left(\frac{1}{\sqrt{2\pi(|\mathcal{Z}|-1)}}\left(\frac{e}{(|\mathcal{Z}|-1)}\left(N+\frac{|\mathcal{Z}|}{2}\right)\right)^{(|\mathcal{Z}|-1)}\right)$
$=(|\mathcal{Z}|-1)\Bigg(1+\log\bigg(\frac{N+\frac{|\mathcal{Z}|}{2}}{|\mathcal{Z}|-1}\bigg)\Bigg)-\frac{\log\bigl(2\pi(|\mathcal{Z}|-1)\bigr)}{2}.$
Plugging the result into (3), we derive a more explicit generalization error bound: for every $\eta\in(0,1)$,
$\mathbf{P}\left(\big|\mathrm{gen}(W,S)\big|\geq\eta\right)$
$\leq\frac{2}{\sqrt{2\pi k}}\left(\frac{e}{k}\left(N+\frac{k+1}{2}\right)\right)^{k}\cdot\exp\left(-\frac{N\eta^{2}}{2\sigma^{2}}\right).$
IV-B Differentially Private Algorithms
We then consider $\varepsilon$-differentially private algorithms. In the regime of $\varepsilon\leq 1$, we derive bounds that are tighter than the general Theorem 8; otherwise, Theorem 8 is better. This phenomenon has also been observed in [1].
Theorem 9.
Let $S=(Z_{1},\ldots,Z_{N})$ be a dataset of size $N$ where $Z_{i}\in\mathcal{Z}$ are sampled i.i.d. from $P_{Z}$. Let $W\in\mathcal{W}$ be a hypothesis obtained from an algorithm $\mathcal{A}$ whose induced distribution is $P_{W|S}$. For all $s\in\mathcal{S}$, if $\mathcal{A}$ is $\varepsilon$-DP,
  • When $\varepsilon\leq 1/N$,
    $\mathcal{L}(S\rightarrow W)\leq(|\mathcal{Z}|-1)N\varepsilon.$
  • When $1/N<\varepsilon\leq 1$,
    $\mathcal{L}(S\rightarrow W)\leq(|\mathcal{Z}|-1)\log\big(e(N\varepsilon+1)\big).$
The proof of Theorem 9 can be found in Appendix H. The proof relies on a special choice of representative sets. This approach can be combined with the refined constructions of representative sets developed in the proofs of Theorems 5 and 6. Consequently, it is possible to integrate these constructions into the current proof of Theorem 9 and further strengthen our maximal leakage bounds. We leave this direction for future work.
V Concluding Remarks
We have provided two classes of explicit generalization error bounds, one based on mutual information and another based on maximal leakage. For the former class, we strictly improve the bounds in [1], while for the latter class, we propose novel bounds. It is nontrivial but of interest to apply the techniques in Section III to the maximal leakage bounds in Section IV. It is possible to improve bounds in both classes through better choices of the mixture distribution in Lemma 2.
Moreover, in [1], another class of private algorithms based on Gaussian differential privacy [44] has been studied. We will also study them in the future, together with other extensions of differential privacy [42, 43].
Appendix A Proof of Claim 2
Proof.
Denote $k:=|\mathcal{Z}|-1$, we calculate
$|\mathcal{T}_{\mathcal{Z},N}|$ $=\binom{N+k}{k}$
$=\frac{(N+k)!}{N!\,k!}$
$=\frac{\prod_{i=1}^{k}(N+i)}{k!}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\left(\frac{1}{k}\sum_{i=1}^{k}(N+i)\right)^{k}}{k!}$
$=\frac{\left(N+\frac{k+1}{2}\right)^{k}}{k!}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{\left(N+\frac{k+1}{2}\right)^{k}}{\sqrt{2\pi k}\,\left(\frac{k}{e}\right)^{k}}$
$=\frac{1}{\sqrt{2\pi k}}\left(\frac{e}{k}\left(N+\frac{k+1}{2}\right)\right)^{k},$
where $(a)$ is by the AM-GM inequality applied to $N+1,\ldots,N+k$ and get $\prod^{k}_{i=1}(N+i)\leq(N+(k+1)/2)^{k}$, and $(b)$ is by the Stirling’s lower bound $k!\geq\sqrt{2\pi k}(k/e)^{k}$. ∎
Appendix B Proof of Lemma 3
Proof.
Since for every measurable set $E$ we have
$Q_{b}(E)\geq\alpha_{b,i}Q_{i}(E),$
it follows that
$Q(E)$ $=\sum_{b=1}^{M}\omega_{b}Q_{b}(E)$
$\geq\sum_{b=1}^{M}\omega_{b}\alpha_{b,i}Q_{i}(E)$
$=:c_{i}\,Q_{i}(E),$
where $c_{i}:=\sum_{b=1}^{M}\omega_{b}\alpha_{b,i}$. Therefore, we have $\log\frac{\mathrm{d}Q_{i}}{\mathrm{d}Q}\leq-\log c_{i}$, $Q_{i}$-a.e.. Since $P\ll Q_{i}$, we obtain
$D_{\mathrm{KL}}(P\|Q)$ $=\mathbf{E}_{P}\left[\log\frac{\mathrm{d}P}{\mathrm{d}Q}\right]$
$=\mathbf{E}_{P}\left[\log\frac{\mathrm{d}P}{\mathrm{d}Q_{i}}+\log\frac{\mathrm{d}Q_{i}}{\mathrm{d}Q}\right]$
$\leq\mathbf{E}_{P}\left[\log\frac{\mathrm{d}P}{\mathrm{d}Q_{i}}\right]-\log c_{i}$
$=D_{\mathrm{KL}}(P\|Q_{i})-\log c_{i},$
which proves the claim. ∎
Appendix C Proof of Theorem 4
Proof.
Fix $t\in\{1,2,\dots,N\}$. Denote the cardinality of alphabet $\mathcal{Z}$ by $m:=|\mathcal{Z}|$ where $m\geq 2$.
Construct the representative set $\mathcal{S}_{0}=\{s_{1},\dots,s_{M}\}$ with parameter $t$ such that $1\leq t\leq N$. Define the uniform mixture
$Q_{W}^{(t)}:=\frac{1}{M}\sum_{b=1}^{M}P_{W|S=s_{b}}.$
Fix an arbitrary dataset $s$. By the same covering argument used for [1, Eq. (52)], there exists some representative $s_{i}\in\mathcal{S}_{0}$ such that
$\mathsf{d}(s,s_{i})\ \leq\ \frac{N}{t}(m-1).$ (12)
Since $\mathcal{A}$ is $\varepsilon$-DP, group privacy implies for all measurable $E$,
$P_{W|S=s_{b}}(E)\ \geq\ e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\,P_{W|S=s_{i}}(E),$
we apply our Lemma 3 with $\omega_{b}=1/M$ and $\alpha_{b,i}=e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$, and get
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}^{(t)}\bigr)$
$\leq D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,P_{W|S=s_{i}}\bigr)-\log\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\Big).$ (13)
Take $k:=\mathsf{d}(s,s_{i})$, by (4),
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,P_{W|S=s_{i}}\bigr)\ \leq\ k\varepsilon.$
Combining with (12) gives
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,P_{W|S=s_{i}}\bigr)\ \leq\ (m-1)\frac{\varepsilon N}{t}.$ (14)
For the case of $t=1$, there is only one grid cube, therefore $M=1$ and $(1/M)\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}=1$, and hence the last term in (13) is 0.
For the case of $t\geq 2$, we have $M\geq 2$ and there exists some $j\neq i$ with $s_{j}\in\mathcal{S}_{0}$. Since for any two size-$N$ datasets $u,v$ one always has $\mathsf{d}(u,v)\leq N$. Therefore,
$\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$
$\geq e^{-\varepsilon\mathsf{d}(s_{i},s_{i})}+e^{-\varepsilon\mathsf{d}(s_{j},s_{i})}$
$\geq 1+e^{-\varepsilon N},$
and hence
$\log\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\Big)\leq\log M-\log\bigl(1+e^{-\varepsilon N}\bigr).$ (15)
Since there are at most $t^{|\mathcal{Z}|-1}$ smaller hypercubes, we have $M\leq t^{m-1}$, and hence $\log M\leq(m-1)\log t$.
Combining (13), (14), and (15), we know for every $s$,
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}^{(t)}\bigr)$
$\leq(m-1)\frac{\varepsilon N}{t}+(m-1)\log t-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon N}\bigr).$
One can choose $t$ minimizing the right-hand side and define $Q_{W}:=Q_{W}^{(t)}$ for that minimizing $t$. ∎
Appendix D Proof of Theorem 5
Proof.
Fix $t\in\{1,2,\dots,N\}$. Let $\mathcal{S}_{0}=\{s_{1},\dots,s_{M}\}$ be the representative set constructed as follows: instead of constructing a covering of the space $0,N]^{|\mathcal{Z}|-1}$ with $t^{|\mathcal{Z}|-1}$ hypercubes, here we only count the hypercubes strictly needed to cover the hypervolume under the $(|\mathcal{Z}|-2)$-simplex, which has $|\mathcal{Z}|-1$ perpendicular edges of length $N$ which intersect at the origin. By [1, Lemma 3], the number of hypercubes needed to cover the hypervolume under the $(|\mathcal{Z}|-2)$-simplex is given by (21). From such $\mathcal{S}_{0}$, we define the uniform mixture
$Q_{W}^{(t)}:=\frac{1}{M}\sum_{b=1}^{M}P_{W|S=s_{b}}.$
From the same covering argument as in the proof of Theorem 4, there exists an $s_{i}\in\mathcal{S}_{0}$ such that
$\mathsf{d}(s,s_{i})\leq(|\mathcal{Z}|-1)\frac{N}{t}.$ (16)
Since $\mathcal{A}$ is $\varepsilon$-DP, for all measurable $E$,group privacy gives,
$P_{W|S=s_{b}}(E)\geq e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}P_{W|S=s_{i}}(E).$
We apply Lemma 3 with $\omega_{b}=1/M$ and $\alpha_{b,i}=e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$, which yields
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}^{(t)}\bigr)$
$\leq D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,P_{W|S=s_{i}}\bigr)-\log\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\Big).$ (17)
Take $k:=\mathsf{d}(s,s_{i})$. By [1, Claim 2] we know
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,P_{W|S=s_{i}}\bigr)\leq k\varepsilon.$
Combining with (16) gives
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,P_{W|S=s_{i}}\bigr)\leq(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}.$ (18)
If $t=1$, then $M=1$ and the overlap term in (17) is 0.
If $t\geq 2$, we have $M\geq 2$ and moreover there exists $j\neq i$ such that
$\mathsf{d}(s_{i},s_{j})\leq\Delta_{t}:=\Big\lceil\frac{N}{t}\Big\rceil+1.$ (19)
Therefore
$\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$ $\geq e^{-\varepsilon\mathsf{d}(s_{i},s_{i})}+e^{-\varepsilon\mathsf{d}(s_{j},s_{i})}$
$\geq 1+e^{-\varepsilon\Delta_{t}},$
and hence
$\log\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\Big)\leq\log M-\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr).$ (20)
Then by the construction of the representative set described at the beginning of the proof (also see [1, Lemma 3]), the number of kept hypercubes satisfies $M\leq S_{|\mathcal{Z}|-1}(t)$ where
$S_{|\mathcal{Z}|-1}(t):=\frac{1}{(|\mathcal{Z}|-1)!}\frac{(t+|\mathcal{Z}|-2)!}{(t-1)!}.$
We have $\log M\leq\log S_{|\mathcal{Z}|-1}(t)$ and combining (17), (18), and (20) gives
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}^{(t)}\bigr)$
$\leq(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+\log S_{|\mathcal{Z}|-1}(t)-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr)$
where
$\Delta_{t}:=\Big\lceil\frac{N}{t}\Big\rceil+1.$
Finally, by [1, Lemma 3],
$S_{|\mathcal{Z}|-1}(t)\leq\frac{1}{(|\mathcal{Z}|-1)!}\left(t+\frac{|\mathcal{Z}|-2}{2}\right)^{|\mathcal{Z}|-1},$ (21)
we derive
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}^{(t)}\bigr)$
$\leq(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+(|\mathcal{Z}|-1)\log\left(t+\frac{|\mathcal{Z}|-2}{2}\right)$
$\qquad-\log(|\mathcal{Z}|-1)!-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr).$
Appendix E Proof of Theorem 6
Proof.
The main idea of the proof is to utilize analyses based on strong typicality, as used in [1, Appendix E], together with Lemma 3.
Recall the definition of strong typical set:
$\mathcal{T}_{\varepsilon}^{N}(Z):=\{Z^{N}:$ $\,\,\big|T_{Z^{N}}(a)-P_{Z}(a)\big|\leq\varepsilon\text{ if }P_{Z}(a)>0,$
$\mathsf{N}(a|Z^{N})=0\text{ otherwise}\},$
and when the context is clear we use $\mathcal{T}$ for short.
By the union bound and Hoeffding’s inequality, we know
$P_{Z}^{\otimes N}(\mathcal{Z}\backslash\mathcal{T})\leq 2|\mathcal{Z}|\exp(-2N\varepsilon^{2}).$ (22)
Take $m:=|\mathcal{Z}|\geq 2$ for the ease of notation.
Take $\Delta_{t}:=\lceil\ell_{t}\rceil+1$ where
$\ell_{t}:=\frac{2\sqrt{N\log N}}{t}.$
Take $\eta:=\sqrt{(\log N)/N}$.
By [1, Eq. (77)],
$I(S;W)$ $\leq\mathbb{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\in\mathcal{T}\right]\mathbb{P}(S\in\mathcal{T})$
$\,\,\,\,\,+\mathbb{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\notin\mathcal{T}\right]\mathbb{P}(S\notin\mathcal{T}).$ (23)
By [1, Eq. (72)], the typical count vector $\mathsf{N}_{s}$ lies inside an $m$-dimensional hypercube of side at most $2\sqrt{N\log N}$. We split each coordinate interval into $t$ parts, giving small hypercubes of side $\ell_{t}\leq 2\sqrt{N\log N}/t$. Take $\mathcal{S}_{0}=\{s_{1},\dots,s_{M}\}\subset\mathcal{T}$ to be the set of representative datasets, and define the uniform mixture
$Q_{W}^{(t)}:=\frac{1}{M}\sum_{b=1}^{M}P_{W|S=s_{b}},$
and we know the distance between any $s$ belonging to the $i$-th hypercube and its center $s_{i}$ is bounded by
$\mathsf{d}(s,s_{i})\leq\frac{m\sqrt{N\log N}}{t}.$ (24)
Fix $s\in\mathcal{T}$ and let $s_{i}$ be the representative of its small hypercube, by group privacy
$D_{\mathrm{KL}}(P_{W|S=s}\|P_{W|S=s_{i}})\leq\varepsilon\,\mathsf{d}(s,s_{i}).$
By our Lemma 3 with $\omega_{b}=1/M$,
$D_{\mathrm{KL}}(P_{W|S=s}\|Q_{W}^{(t)})$
$\leq\varepsilon\,\mathsf{d}(s,s_{i})-\log\!\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\,\mathsf{d}(s_{b},s_{i})}\Big).$ (25)
If $t=1$, then $M=1$ and the logarithmic term in (25) equals 0.
If $t\geq 2$, since the small hypercubes form a regular grid and we keep only hypercubes that contain at least one type with total count $N$, the kept set satisfies $M\geq 2$ for $t\geq 2$ in the typical region described above. Hence, for each fixed $i$ we can pick some $j\neq i$ and trivially bound $\mathsf{d}(s_{i},s_{j})\leq\Delta_{t}$ with
$\Delta_{t}:=\lceil\ell_{t}\rceil+1\quad\text{and}\quad\ell_{t}=\frac{2\sqrt{N\log N}}{t},$
which yields
$\log\!\Big(\frac{1}{M}\sum_{b=1}^{M}e^{-\varepsilon\,\mathsf{d}(s_{b},s_{i})}\Big)\leq\log M-\log(1+e^{-\varepsilon\Delta_{t}}).$
It remains to upper bound $M$.
At first, since the $m$-dimensional hypercube is partitioned into $t^{m}$ smaller hypercubes, we have $M\leq t^{m}$ and hence
$\log M\leq m\log t.$ (26)
For each choice of the first (m1)(m-1) coordinate bins, the constraint $\sum_{a\in\mathcal{Z}}\mathsf{N}(a|s)=N$ restricts the last coordinate to a range of length at most $(m-1)\ell_{t}$, so it can intersect at most $m$ bins of length $\ell_{t}$. Hence the number of kept hypercubes satisfies $M\leq m\,t^{m-1}$ and hence
$\log M\leq\log(m\,t^{m-1}).$ (27)
Combining (26) and (27), we have for $t\geq 2$,
$\log M\leq\min\{\,m\log t,\ \log(m\,t^{m-1})\,\}.$
Therefore, combining (24) and (25) gives, for all $s\in\mathcal{T}$,
$D_{\mathrm{KL}}(P_{W|S=s}\|Q_{W}^{(t)})$
$\leq\frac{\varepsilon m\sqrt{N\log N}}{t}-\mathds{1}_{\{t\geq 2\}}\log(1+e^{-\varepsilon\Delta_{t}}),$
$\qquad+\mathds{1}_{\{t\geq 2\}}\min\{|\mathcal{Z}|\log t,\log(|\mathcal{Z}|t^{|\mathcal{Z}|-1})\}.$
For the first term in (23), taking conditional expectation over $S\in\mathcal{T}$ gives the same bound.
For the second term, [1, Lemma 2] and Jensen’s inequality give $D_{\mathrm{KL}}(P_{W|S=s}\|Q_{W}^{(t)})\leq\varepsilon N$ for $s\notin\mathcal{T}$. Multiplied by $\mathbb{P}(S\notin\mathcal{T})\leq 2mN^{-2}$ to obtain
$\mathbb{E}\!\left[D_{\mathrm{KL}}(P_{W|S}\|Q_{W}^{(t)})\,\middle|\,S\notin\mathcal{T}\right]\mathbb{P}(S\notin\mathcal{T})\leq\frac{2m\varepsilon}{N}.$
Substituting the typical and atypical bounds into (23) yields (11). ∎
Appendix F Proof of Lemma 7
Proof.
For each $w$, choose an index $i_{w}$ that maximizes $P_{i}(w)$ (we break ties arbitrarily). Then
$\sum_{w}\max_{i}P_{i}(w)$ $=\sum_{w}P_{i_{w}}(w)$
$=\sum^{M}_{i=1}\sum_{w:\,i_{w}=i}P_{i}(w)$
$\leq\sum^{M}_{i=1}\sum_{w}P_{i}(w)$
$=M.$
It remains to characterize equality. Note that for each fixed $w$, we always have
$\max_{1\leq i\leq M}P_{i}(w)\ \leq\ \sum_{i=1}^{M}P_{i}(w),$
with equality if and only if at most one of the numbers $\{P_{i}(w)\}_{i=1}^{M}$ is nonzero. Therefore,
$\sum_{w\in\mathcal{W}}\max_{i}P_{i}(w)=\sum_{w\in\mathcal{W}}\sum_{i=1}^{M}P_{i}(w)$
holds if and only if for every $w\in\mathcal{W}$, there exists at most one index $i$ such that $P_{i}(w)>0$. ∎
Appendix G Proof of Theorem 8
Proof.
Note we assume the algorithm $\mathcal{A}$ is permutation invariant, and therefore $\mathcal{L}(S\rightarrow W)=\mathcal{L}(T_{S}\rightarrow W)$ where $T_{S}$ is the type of $S$. This can be proved as follows: due to the permutation invariance, we know $P_{W|S=s}=P_{W|S=s^{\prime}}$ if $T_{s}=T_{s^{\prime}}$. Hence there exists a conditional distribution $P_{W|T}$ such that for every dataset $s$, $P_{W|S=s}(\cdot)=P_{W|T}(\cdot\,|\,T_{s})$, and
$\mathcal{L}(S\rightarrow W)$ $=\log\sum_{w}\max_{s\in\mathcal{S}:P_{Z}(s)>0}P_{W|S}(w|s)$
$=\log\sum_{w}\max_{s\in\mathcal{S}:P_{Z}(s)>0}P_{W|T}(w|T_{s})$
$=\log\sum_{w}\max_{t:P_{T_{S}}(t)>0}P_{W|T}(w|t)$
$=\mathcal{L}(T_{S}\rightarrow W)$
$\leq\log|\mathrm{supp}(T_{S})|$
$\leq\log|\mathcal{T}_{\mathcal{Z},N}|,$
where the inequality $\mathcal{L}(T_{S}\rightarrow W)\leq\log|\mathrm{supp}(T_{S})|$ follows by applying Lemma 7 to the family of pmfs $\{P_{W|T=t}\}_{t\in\mathrm{supp}(T_{S})}$, and the last inequality holds since $T_{S}\in\mathcal{T}_{\mathcal{Z},N}$. Then because $|\mathcal{T}_{\mathcal{Z},N}|\leq(N+1)^{|\mathcal{Z}|-1}$ [1, Claim 1], we directly obtain
$\mathcal{L}(S\rightarrow W)\leq\left(|\mathcal{Z}|-1\right)\log(N+1).$
Appendix H Proof of Theorem 9
Proof.
Fix $t\in\{1,\ldots,N\}$. Similar to [1] we partition the $(|\mathcal{Z}|-1)$ dimensional simplex by hypercubes of side $1/t$. For each non-empty cell (which contains at least one type $T_{s}$), we choose one dataset $s_{0}$ whose type lies in that cell and collect all such representatives in the set $\mathcal{S}_{0}$. It is straightforward to check that
$|\mathcal{S}_{0}|\leq t^{|\mathcal{Z}|-1}.$
Note if $T_{s}$ and $T_{s_{0}}$ are in the same cell, then for coordinates $a=1,\ldots,|\mathcal{Z}|-1$ we know
$|T_{s}(a)-T_{s_{0}}(a)|\leq\frac{1}{t}$
and also the last coordinate satisfies
$|T_{s}(|\mathcal{Z}|)-T_{s_{0}}(|\mathcal{Z}|)|$ $=\left|\sum^{|\mathcal{Z}|-1}_{a=1}\big(T_{s}(a)-T_{s_{0}}(a)\big)\right|$
$\leq\sum^{|\mathcal{Z}|-1}_{a=1}\left|T_{s}(a)-T_{s_{0}}(a)\right|$
$\leq\frac{|\mathcal{Z}|-1}{t}.$
Therefore
$\|T_{s}-T_{s_{0}}\|_{1}$ $\leq\sum^{|\mathcal{Z}|}_{a=1}\left|T_{s}(a)-T_{s_{0}}(a)\right|$
$\leq\frac{2(|\mathcal{Z}|-1)}{t}.$
Note we use the multiset distance [1]:
$\mathsf{d}(s,s_{0})$ $=\frac{1}{2}\sum_{a\in\mathcal{Z}}\left|\mathsf{N}(a|s)-\mathsf{N}(a|s_{0})\right|$
$=\frac{N}{2}\|T_{s}-T_{s_{0}}\|_{1}$
$\leq\frac{N}{2}\cdot\frac{2(|\mathcal{Z}|-1)}{t}$
$=\frac{N}{t}(|\mathcal{Z}|-1).$
For each $s\in\mathcal{S}$, let $s_{0}(s)\in\mathcal{S}_{0}$ be the representative whose type lies in the same cell as $T_{s}$, and define $k(s):=d(s,s_{0}(s))$. By the covering argument above, for all $s$,
$k(s)\leq K:=\frac{N}{t}(|\mathcal{Z}|-1).$
Hence, for all $s\in\mathcal{S}$ and all $w\in\mathcal{W}$,
$P_{W|S=s}(w)$ $\leq e^{\varepsilon k(s)}P_{W|S=s_{0}(s)}(w)$
$\leq e^{\varepsilon k(s)}\max_{\bar{s}\in\mathcal{S}_{0}}P_{W|S=\bar{s}}(w)$
$\leq e^{\varepsilon K}\max_{\bar{s}\in\mathcal{S}_{0}}P_{W|S=\bar{s}}(w).$
Taking the maximum over $s$ yields
$\max_{s\in\mathcal{S}}P_{W|S=s}(w)\leq e^{\varepsilon K}\max_{\bar{s}\in\mathcal{S}_{0}}P_{W|S=\bar{s}}(w).$
Summing over $w$ and invoking Lemma 7 on $\{P_{W|S=\bar{s}}\}_{\bar{s}\in\mathcal{S}_{0}}$, we obtain
$\sum_{w\in\mathcal{W}}\max_{s\in\mathcal{S}}P_{W|S=s}(w)$ $\leq e^{\varepsilon K}\sum_{w\in\mathcal{W}}\max_{\bar{s}\in\mathcal{S}_{0}}P_{W|S=\bar{s}}(w)$
$\leq e^{\varepsilon K}\,|\mathcal{S}_{0}|.$
Therefore, by the definition of maximal leakage,
$\mathcal{L}(S\to W)$ $=\log\sum_{w\in\mathcal{W}}\max_{s\in\mathcal{S}}P_{W|S=s}(w)$
$\leq\varepsilon K+\log|\mathcal{S}_{0}|$
$\leq\varepsilon\frac{N}{t}(|\mathcal{Z}|-1)+(|\mathcal{Z}|-1)\log t$
$=(|\mathcal{Z}|-1)\left(\frac{N\varepsilon}{t}+\log t\right).$
We then optimize over $t$, and obtain the following result:
  • When $\varepsilon\leq 1/N$,
    $\mathcal{L}(S\rightarrow W)\leq(|\mathcal{Z}|-1)N\varepsilon.$
  • When $1/N<\varepsilon\leq 1$,
    $\mathcal{L}(S\rightarrow W)\leq(|\mathcal{Z}|-1)\log\big(e(N\varepsilon+1)\big).$